OpenStudy (anonymous):

need help with x/4+7=-2

OpenStudy (ddcamp):

(x/4) + 7 or x/(4+7)?

OpenStudy (anonymous):

need to understand how to find what X equals

OpenStudy (ddcamp):

Yes, I understand that. But I need to know whether or not the 7 is under the fraction bar. Assuming it isn't, start by adding 7 to both sides to get: x/4 = 5 Then multiply by 4 to solve for x.

OpenStudy (anonymous):

X over 4 +7=-2

OpenStudy (ddcamp):

(x/4)+7 -7 = -2 -7 (x/4) = -9 (x/4)×4 = -9×4 x = -36

OpenStudy (anonymous):

Thank you